# Imperative Learning: A Self-supervised Neuro-Symbolic Learning Framework for Robot Autonomy

### v6

@@ 5.4.2 Metrics
Following prior works (Qiu et al., 2022), we use the widely-used absolute trajectory error (ATE), relative motion error (RME), and root mean square error (RMSE) of rotational and translational drifts as evaluation metrics.
5.4.3 Accuracy Evaluation
We first assess the localization accuracy of iSLAM on the KITTI dataset, which has been widely used in previous works on various sensor setups. To facilitate a fair comparison, in Table 12, we evaluate our standalone VO component against existing VO networks, including DeepVO (Wang et al., 2017b), UnDeepVO (Li et al., 2018), TartanVO (Wang et al., 2021b), and DROID-SLAM (Teed and Deng, 2021), and compare the full iSLAM to other learning-based visual-inertial methods, which are Wei et al. (2021), Yang et al. (2022a), and DeepVIO (Han et al., 2019). Sequences 00 and 03 are omitted in our experiment since they lack completed IMU data. Notably, some methods such as DeepVO and (Yang et al., 2022a) were supervisedly trained on KITTI, while our iSLAM was self-supervised. Even though, iSLAM outperforms all the competitors. Additionally, it is noteworthy that our base model, TartanVO, doesn’t exhibit the highest performance due to its lightweight design. Nevertheless, through imperative learning, we achieve much lower errors with a similar model architecture.
We also test iSLAM on the EuRoC benchmark, where the presence of aggressive motion, substantial IMU drift, and significant illumination changes pose notable challenges to SLAM algorithms (Qin et al., 2018). However, both our standalone VO and the full iSLAM generalize well to EuRoC. As shown in Table 13, iSLAM achieves an average ATE 62% lower than DeepV2D (Teed and Deng, 2018), 76% lower than DeepFactors (Czarnowski et al., 2020), and 42% lower than TartanVO (Wang et al., 2021b).
@@ 5.4.4 Efficiency Analysis
Efficiency is a crucial factor for SLAM systems in real-world robot applications. Here we conduct the efficiency assessments for iSLAM on an RTX4090 GPU. Our stereo VO can reach a real-time speed of 29-31 frames per second (FPS). The scale corrector only uses about 11% of inference time, thus having minimal impact on the overall efficiency. The IMU module achieves an average speed of 260 FPS, whereas the back-end achieves 64 FPS, when evaluated independently. The entire system operates at around 20 FPS. Note that the imperative learning framework offers broad applicability across different front-end and back-end designs, allowing for great flexibility in balancing accuracy and efficiency. We next measure the runtime of our “one-step” back-propagation strategy against the conventional unrolling approaches (Tang and Tan, 2018; Teed and Deng, 2021) during gradient calculations. A significant runtime gap is observed: our “one-step” strategy is on average 1.5$\times$ faster than the unrolling approach.
@@ 5.5.1 Baselines
To thoroughly evaluate the performance of different methods for MTSP, we compare our iMTSP with both classic methods and learning-based methods. This includes a well-known meta-heuristic method implemented by the Google OR-Tools routing library (Furnon and Perron, 2023) and an RL-based method (Hu et al., 2020), which will be referred to as Google OR-Tools and RL baseline for abbreviation, respectively. Specifically, Google OR-Tools can provide near-optimal solutions [^3] Google OR-Tools is a meta-heuristic and cannot guarantee the optimality of solutions. However, it is efficient for small-scale problems and wildly used for academic and commercial purposes. Thus, in our experiments, it can be seen as an “ideal” solver providing “optimal” solutions. to MTSPs with a few hundred cities, while the RL baseline can solve larger-scale MTSPs such as with one thousand cities. We cannot compare with other methods such as (Liang et al., 2023; Gao et al., 2023; Park et al., 2021) because they either don’t provide source code or cannot be applied to our settings.
Table 14: Performance of different methods on large-scale MTSPs. iMTSP significantly outperforms the Google OR-Tools (Furnon and Perron, 2023) and provides an average of $3.2\pm 0.01\%$ shorter maximum route length than the RL baseline (Hu et al., 2020).
@@
Figure 17: The performance visualization of the baselines and iMTSP for a problem instance with a central depot, $5$ agents, and $500$ cities. The numbers denote the length of the route, and different colors represent different routes. The longest tour lengths are underlined and iMTSP has the best solution and has fewer sub-optimal patterns like circular partial routes or long straight routes.
@@ 5.5.2 Datasets
Due to the uniqueness of MTSP, we cannot deploy MTSP algorithms on such a large-scale robot team, thus we build a simulation for comparison. Specifically, all the data instances are generated by uniformly sampling points in a unit rectangular so that both $x$ and $y$ coordinates are in the range of 00 to $1$. The test set consists of i.i.d. samples but with a larger number ($400$ to $1000$) of cities as aforementioned. Note that the proposed approach is applicable to general graphs with arbitrary costs assigned to edges, although our dataset is constrained in the 2-D space.
@@ 5.5.3 Implementation Details
The structure of our allocation network is similar to that in (Hu et al., 2020), where the cities and agents are embedded into $64$ dimensional vectors. The surrogate network is a three-layer MLP whose hidden dimension is $256$ and whose activation function is “tanh”. Additionally, Google OR-Tools provides the TSP solver in our iMTSP. It uses the “Global Cheapest Arc” strategy to generate its initial solutions and uses “Guided local search” for local improvements. We use the best-recommended settings in Google OR Tools for all experiments. All tests were conducted locally using the same desktop-level computer.
@@ 5.5.4 Quantitative Performance
Table 15: Computing time of iMTSP dealing with $1000$ cities. Due to a similar network structure, the RL baseline (Hu et al., 2020) has roughly the same runtime efficiency compared with ours, while Google OR-Tools (Furnon and Perron, 2023) cannot find feasible solutions within a given time budget.
@@ 5.5.5 Efficiency Analysis
Figure 18: The gradient variance history of our method and the RL baseline during the training process with $50$ and $100$ cities. The vertical axis is the logarithmic variance of the allocation network gradient, and the horizontal axis denotes the number of training iterations. Our control-variate-base gradient estimator converges about 20×20$\times$20 × faster than the RL baseline.
@@ 5.5.7 Gradient Variance
As demonstrated in Section 4.5, our control variate-based iMTSP framework is expected to have a smaller gradient variance than the RL baseline. We verify such a hypothesis by explicitly recording the mini-batch gradient variance during the training process. The experiment is conducted twice with $10$ agents respectively visiting $50$ and $100$ cities. The training data with a batch size of $512$ is divided into several mini-batches containing $32$ instances. We compute and store the gradients for every mini-batch, and then calculate their variance for the entire batch. The parameters of the allocation network are still updated with the average gradient of the whole batch. The variance of the gradient with respect to the training process is shown in Figure 18, where the horizontal axis represents the number of training iterations and the vertical axis denotes the mean logarithm gradient variance. It is worth noting that the gradient of our iMTSP converges about 20×20$\times$20 × faster than the RL baseline (Hu et al., 2020), which verifies the effectiveness of our control variate-based BLO process.
@@ 6 Conclusions & Discussions
We believe that the theoretical challenges of imperative learning will also inspire new directions and topics for fundamental research in bilevel optimization. For example, second-order bilevel optimization remains underexplored in machine learning due to time and memory constraints. However, its relevance is increasingly critical in robotics, where second-order optimization is essential for achieving the required accuracy in complex tasks. Furthermore, handling lower-level constraints within general bilevel optimization settings presents significant challenges and remains an underdeveloped area. Recent advancements, such as those based on proximal Lagrangian value functions (Yao et al., 2024a, 2023), offer potential solutions for constrained robot learning problems. We intend to investigate robust approaches and apply these techniques to enhance our experimental outcomes for imperative learning. Additionally, we plan to develop heuristic yet practical solutions for bilevel optimization involving discrete variables, and we aim to refine the theoretical framework for these discrete scenarios with fewer assumptions, leveraging recent progress in control variate estimation.
To enhance the usability of imperative learning for robot autonomy, we will extend PyPose (Wang et al., 2023; Zhan et al., 2023), an open-source Python library for robot learning, to incorporate bilevel optimization frameworks. Additionally, we will provide concrete examples of imperative learning using PyPose across various domains, thereby accelerating the advancement of robot autonomy.

### v7

@@ 5.4.2 Metrics
Following prior works (qiu2021airdos), we use the widely-used absolute trajectory error (ATE), relative motion error (RME), and root mean square error (RMSE) of rotational and translational drifts as evaluation metrics.
@@
5.4.3 Accuracy Evaluation
We first assess the localization accuracy of iSLAM on the KITTI dataset, which has been widely used in previous works on various sensor setups. To facilitate a fair comparison, in Table 12, we evaluate our standalone VO component against existing VO networks, including DeepVO (supe_wang2017deepvo), UnDeepVO (unsup_li2018undeepvo), TartanVO (wang2021tartanvo), and DROID-SLAM (teed2021droid), and compare the full iSLAM to other learning-based visual-inertial methods, which are wei2021unsupervised, yang2022efficient, and DeepVIO (han2019deepvio). Sequences 00 and 03 are omitted in our experiment since they lack completed IMU data. Notably, some methods such as DeepVO and (yang2022efficient) were supervisedly trained on KITTI, while our iSLAM was self-supervised. Even though, iSLAM outperforms all the competitors. Additionally, it is noteworthy that our base model, TartanVO, doesn’t exhibit the highest performance due to its lightweight design. Nevertheless, through imperative learning, we achieve much lower errors with a similar model architecture.
We also test iSLAM on the EuRoC benchmark, where the presence of aggressive motion, substantial IMU drift, and significant illumination changes pose notable challenges to SLAM algorithms (qin2018vins). However, both our standalone VO and the full iSLAM generalize well to EuRoC. As shown in Table 13, iSLAM achieves an average ATE 62% lower than DeepV2D (supe_teed2018deepv2d), 76% lower than DeepFactors (czarnowski2020deepfactors), and 42% lower than TartanVO (wang2021tartanvo).
@@ 5.4.4 Efficiency Analysis
[…]
@@ 5.5.1 Baselines
To thoroughly evaluate the performance of different methods for MTSP, we compare our iMTSP with both classic methods and learning-based methods. This includes a well-known meta-heuristic method implemented by the Google OR-Tools routing library (ortools_routing) and an RL-based method (hu2020reinforcement), which will be referred to as Google OR-Tools and RL baseline for abbreviation, respectively. Specifically, Google OR-Tools can provide near-optimal solutions [^3] Google OR-Tools is a meta-heuristic and cannot guarantee the optimality of solutions. However, it is efficient for small-scale problems and wildly used for academic and commercial purposes. Thus, in our experiments, it can be seen as an “ideal” solver providing “optimal” solutions. to MTSPs with a few hundred cities, while the RL baseline can solve larger-scale MTSPs such as with one thousand cities. We cannot compare with other methods such as (liang2023splitnet; gao2023amarl; park2021schedulenet) because they either don’t provide source code or cannot be applied to our settings.
Table 14: Performance of different methods on large-scale MTSPs. iMTSP significantly outperforms the Google OR-Tools (ortools_routing) and provides an average of $3.2\pm 0.01\%$ shorter maximum route length than the RL baseline (hu2020reinforcement).
@@
Figure 17: The performance visualization of the baselines and iMTSP for a problem instance with a central depot, 55 agents, and $500$ cities. The numbers denote the length of the route, and different colors represent different routes. The longest tour lengths are underlined and iMTSP has the best solution and has fewer sub-optimal patterns like circular partial routes or long straight routes.
@@ 5.5.2 Datasets
Due to the uniqueness of MTSP, we cannot deploy MTSP algorithms on such a large-scale robot team, thus we build a simulation for comparison. Specifically, all the data instances are generated by uniformly sampling points in a unit rectangular so that both $x$ and $y$ coordinates are in the range of 0 to $1$. The test set consists of i.i.d. samples but with a larger number ($400$ to $1000$) of cities as aforementioned. Note that the proposed approach is applicable to general graphs with arbitrary costs assigned to edges, although our dataset is constrained in the 2-D space.
@@ 5.5.3 Implementation Details
The structure of our allocation network is similar to that in (hu2020reinforcement), where the cities and agents are embedded into $64$ dimensional vectors. The surrogate network is a three-layer MLP whose hidden dimension is $256$ and whose activation function is “tanh”. Additionally, Google OR-Tools provides the TSP solver in our iMTSP. It uses the “Global Cheapest Arc” strategy to generate its initial solutions and uses “Guided local search” for local improvements. We use the best-recommended settings in Google OR Tools for all experiments. All tests were conducted locally using the same desktop-level computer.
@@ 5.5.4 Quantitative Performance
Table 15: Computing time of iMTSP dealing with $1000$ cities. Due to a similar network structure, the RL baseline (hu2020reinforcement) has roughly the same runtime efficiency compared with ours, while Google OR-Tools (ortools_routing) cannot find feasible solutions within a given time budget.
@@ 5.5.5 Efficiency Analysis
Figure 18: The gradient variance history of our method and the RL baseline during the training process with $50$ and $100$ cities. The vertical axis is the logarithmic variance of the allocation network gradient, and the horizontal axis denotes the number of training iterations. Our control-variate-base gradient estimator converges about $20\times$ faster than the RL baseline.
@@ 5.5.7 Gradient Variance
As demonstrated in Section 4.5, our control variate-based iMTSP framework is expected to have a smaller gradient variance than the RL baseline. We verify such a hypothesis by explicitly recording the mini-batch gradient variance during the training process. The experiment is conducted twice with $10$ agents respectively visiting $50$ and $100$ cities. The training data with a batch size of $512$ is divided into several mini-batches containing $32$ instances. We compute and store the gradients for every mini-batch, and then calculate their variance for the entire batch. The parameters of the allocation network are still updated with the average gradient of the whole batch. The variance of the gradient with respect to the training process is shown in Figure 18, where the horizontal axis represents the number of training iterations and the vertical axis denotes the mean logarithm gradient variance. It is worth noting that the gradient of our iMTSP converges about $20\times$ faster than the RL baseline (hu2020reinforcement), which verifies the effectiveness of our control variate-based BLO process.
@@ 6 Conclusions & Discussions
We believe that the theoretical challenges of imperative learning will also inspire new directions and topics for fundamental research in bilevel optimization. For example, second-order bilevel optimization remains underexplored in machine learning due to time and memory constraints. However, its relevance is increasingly critical in robotics, where second-order optimization is essential for achieving the required accuracy in complex tasks. Furthermore, handling lower-level constraints within general bilevel optimization settings presents significant challenges and remains an underdeveloped area. Recent advancements, such as those based on proximal Lagrangian value functions (yao2024overcoming; yao2023constrained), offer potential solutions for constrained robot learning problems. We intend to investigate robust approaches and apply these techniques to enhance our experimental outcomes for imperative learning. Additionally, we plan to develop heuristic yet practical solutions for bilevel optimization involving discrete variables, and we aim to refine the theoretical framework for these discrete scenarios with fewer assumptions, leveraging recent progress in control variate estimation.
To enhance the usability of imperative learning for robot autonomy, we will extend PyPose (wang2023pypose; zhan2023pypose), an open-source Python library for robot learning, to incorporate bilevel optimization frameworks. Additionally, we will provide concrete examples of imperative learning using PyPose across various domains, thereby accelerating the advancement of robot autonomy.